Circular No. 8337

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2004 HV_{60} (SPACEWATCH) An apparently asteroidal object discovered by the Spacewatch survey

(discovery observation below from MPS 106357 and MPEC 2004-H91) has been reported by D. J. Tholen, University of Hawaii (UH), to be cometary on two 400-s R exposures taken on May 8.3 UT by F. Bernardi, J. Pittichova, and N. Moskovitz with the UH 2.24-m telescope on Mauna Kea. There is a faint tail ≈ 55′′ long in p.a. ∼ 24^{o} and a hint of a second, fainter tail almost 15′′ long in p.a. 310^{o}. The measured PSF widths of the comet head are 1′′.03 and 0′′.95 in the two images, while the widths of 55 trailed astrometric reference stars in the images range from 0′′.90 to 1′′.09, not sufficient to resolve the presence of a coma. R. H. McNaught reports that images obtained on Apr. 30.6 at Siding Spring also suggest a slightly soft image for the comet.

2004 UT  α_{2000}  δ_{2000}  Mag.
Apr. 25.34519  15^{h}06^{m}21.^{s}20  −16^{o}58′40′′.9  21.4

Additional astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2004-J34.

T = 2003 Dec. 21.526 TT  ω = 149.^{o}127
Ω = 44.849 }2000.0
q = 3.10499 AU  i = 93.488

SUPERNOVA 2004bl IN MCG +00-31-42 G. H. Marion and C. L. Gerardy, University of Texas at Austin, report

that a low-resolution optical spectrogram (resolution 300; range 410–1000 nm) of SN 2004bl (cf. IAUC 8334), obtained on May 8.16 UT with the 9.2m Hobby-Eberly Telescope (+ Marcario Low-Resolution Spectrograph) by S. Odewahn and M. Graver, shows it to be a type-Ia supernova within a few days of maximum light; adopting the NED redshift of 5192 km/s for MCG +00-31-42, the expansion velocity of the Si II feature (rest wavelength 635.5 nm) is found to be 11300 km/s.

COMET C/2002 T7 (LINEAR) Visual total-magnitude estimates: Apr. 29.87 UT, 4.1 (A. Pearce, No-

ble Falls, W. Australia, 8×40 binoculars; 7^{o}.6 tail); May 2.79, 3.8 (D. A. J. Seargent, The Entrance, N.S.W., naked eye); 9.35, 3.6 (A. Amorim, Florianopolis, Brazil, 7×50 binoculars; moonlight).

2004 May 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green